OpenStudy (albert0898):

Please help me understand this. 26. What positive value of k would make the lines below parallel in the (x, y) coordinate plane? kx + 2y = 6 32x + ky = 18 F. 0 G. 3 H. 6 J. 8 K. 12 Thank you. EDIT: Apart from substituting the numbers in, what's another method? This is to be treated as a timed test, so the fastest method possible is key.

OpenStudy (asnaseer):

for the lines to be parallel, the ration between the x and y terms must be the same

OpenStudy (albert0898):

That I understand, but the variable k is confusing me so much

OpenStudy (asnaseer):

that therefore gives us that:\[\frac{32}{k}=\frac{k}{2}\]

OpenStudy (asnaseer):

make sense?

OpenStudy (albert0898):

One sec, let me work this out on paper.

OpenStudy (asnaseer):

i.e. the second line is some multiple of the first line

OpenStudy (albert0898):

Okay, so when I brought the terms over, I got... kx + 2y = 6 becomes y = 3 - kx/2 y = 18/k - 32x/k

OpenStudy (asnaseer):

are you sure you have copied the question down correctly?

OpenStudy (albert0898):

So I bring the slopes down... -kx/2 = -32x/k -64x = -k^2x 64 = k^2 8 = k

OpenStudy (albert0898):

Yes. This is from an ACT practice test

OpenStudy (asnaseer):

OK - you have worked out the correct answer :)

OpenStudy (albert0898):

I couldn't have done it without your assistance. Thank you!

OpenStudy (asnaseer):

yw :)